# BalanceKV: KV Cache Compression through Discrepancy Theory

### v1

@@ 3.2 MergeAndReduce
$\|\varphi(k)\|^{2}_{2}=\exp\left(\frac{\|k\|^{2}_{2}}{\sqrt{d}}\right).$
Consider the set of vectors $\varphi(k_{1})\otimes v_{1},\varphi(k_{2})\otimes v_{2},\ldots$. Run the Self-Balancing Walk algorithm on the set of vectors $\varphi(k_{1})\otimes v_{1},\varphi(k_{2})\otimes v_{2},\ldots$ with failure parameter set to $\delta/s$ and denote by $C^{\prime}$ and $C\backslash C^{\prime}$ the partition of $C$ returned by the algorithm. Observe that, even though vectors $\varphi(k_{i})\otimes v_{i}$ are infinite dimensional, Self-Balancing Walk still can be implemented. The algorithm never has to keep these vectors in the memory because the only operation which requires the knowledge of the embeddings – the inner product – can be performed if we just store vector pairs $\{k_{i},v_{i}\}$:
Denote by $e_{1},\ldots,e_{s}$ the standard orthonormal basis in $\mathbb{R}^{s}$. By Theorem 3.4, for any $i\in[s]$ with probability $1-\delta/s$
and so with probability at least $1-\delta$ all of the above inequalities hold simultaneously.
To simplify the right hand side, notice that $\left\|\varphi(k)\otimes v\right\|_{2}=\exp\left(\frac{\|k\|^{2}_{2}}{2\sqrt{d}}\right)\cdot\|v\|_{2}$ and $\|\varphi(q)\otimes e_{i}\|_{2}=\exp\left(\frac{\|q\|^{2}_{2}}{2\sqrt{d}}\right)$. Observe that for any $i$ it holds that $\langle\varphi(k)\otimes v,\varphi(q)\otimes e_{i}\rangle=\langle\varphi(k),\varphi(q)\rangle\cdot[v]_{i}=\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)\cdot[v]_{i}$ , where we denote the $i$-th coordinate of the vector $v$ by $[v]_{i}$. Therefore, the left hand side of the expression above is simply the absolute value of the $i$-th coordinate of the vector
This completes the proof of Theorem 3.2. ∎
4 Experiments
Experiments in Section 4.1 are performed on a single NVIDIA A100 GPU with 40GB VRAM, and in Section 4.2 on a single NVIDIA RTX A6000 GPU with 48GB VRAM.
4.1 Single Layer Attention Approximation
We first empirically evaluate the performance of BalanceKV for approximating a single attention layer, and compare it with independent uniform sampling. We use the pretrained $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ [DJP+24] and $\mathtt{Ministral}$-$\mathtt{8B}$-$\mathtt{Instruct}$-$\mathtt{2410}$ [Mis24] and TriviaQA dataset from LongBench [BLZ+23], and consider the $1^{st}$,$2^{nd}$ and $5^{th}$ layers of the models for attention approximation.
For given a prompt with length $n$, we store the corresponding query, key, and value embeddings for all layers. Denote a pair of embeddings in some layer by $(q_{1},k_{1},v_{1}),\ldots,(q_{n},k_{n},v_{n})$ and the goal is to approximate the attention $\text{Attn}(q_{j},K_{j},V_{j})$ for the latest $256$ queries, i.e. $j\in[n-256,n]$. Specifically, we keep several first and recent tokens separately and apply BalanceKV to the intermediate row vectors in $K_{j}$. This is motivated by [XTC+23] as important contexts are likely contained in the first and latest tokens. We retain the first 256 embeddings and the recent ones from $n-256$ to $j$ and our compressed cache contains tokens whose indices are in $[256]\cup S\cup\{n-256,\dots,j\}$ where $S\subseteq[257,n-256]$ can be obtained from BalanceKV. We explore four compression parameters $T\in\{1,2,3,4\}$ which reduces the cache memory by a factor of $2^{-T}$. Let $z_{j}$ be our approximation using BalanceKV plus the recent and first few embeddings at the stream $j\in[n-256,n]$. We compute relative errors $\|z_{j}-\text{Attn}(q_{j},K_{j},V_{j})\|_{F}/\|\text{Attn}(q_{j},K_{j},V_{j})\|_{F}$ for all $j\in[n-256,n]$, batches, heads and input prompts in the dataset. We repeat this with 10 different random seeds and compute their average and standard deviations.
We also compare our method to independent uniform sampling, in which we replace the application of BalanceKV with sampling a $2^{-T}$ fraction of key and value embeddings with indices in $[257,n-256]$ uniformly at random. We report the results for layers 1,2 and 5 of $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ and $\mathtt{Ministral}$-$\mathtt{8B}$-$\mathtt{Instruct}$-$\mathtt{2410}$ in Fig. 2. We observe that for each model, layer and compression rate, the error of our method is much lower than independent uniform sampling. This verifies that BalanceKV has a better error bound empirically, confirming the prediction provided by Theorem 3.1.
Figure 2: Comparison of relative errors across different layers of $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ (first figure on top) and $\mathtt{Ministral}$-$\mathtt{8B}$-$\mathtt{Instruct}$-$\mathtt{2410}$ (second figure on the bottom) on TriviaQA dataset.
4.2 End-to-end Evalution on LongBench
We benchmark our algorithm on LongBench [BLZ+23], a comprehensive collection of datasets designed to evaluate the long-context understanding capabilities of large language models. The benchmark consists of various long-text application scenarios, including single-document question-answering, multi-document question-answering, summarization, few-shot learning, synthetic tasks and code completion (Code). In particular, we use the LongBench-E split, which consists of datasets with uniform length distribution.
Our approach compresses the KV cache while generating in the prefill stage but maintains all streamed embeddings $(q_{j},k_{j},v_{j})$ during generation. This is because the number of generated tokens is much smaller than the input sequence length. We set $b=256$ and $T=2$, achieving a consistent compression rate of $0.25$ across all inputs.
We evaluate our method against several token-level KV cache compression schemes, including StreamingLLM [XTC+23], SnapKV [LHY+24], and PyramidKV [CZG+24] as well as uniform sampling described in Section 4.1. We use their implementations written in MInference [JLZ+24] [^1] https://github.com/microsoft/MInference/, and configure their hyperparameters to match a uniform compression rate with $0.25$. We follow the same evaluation metrics from [BLZ+23]. We test them on $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ as before. The results are summarized in Table 1.
Our method achieves the highest average scores among all benchmark approaches. Notably, it outperforms summarization tasks by a large margin, e.g., 23.82 against others except for the uniform sampling. This result justifies that a subset obtained from discrepancy theory has practical impacts on various LLM tasks.
Table 1: Performance evaluation on LongBench tasks with $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ across various token-level KV cache compression methods. A higher score is better.
Method Task Average
Single-Doc QA Multi-Doc QA Summarization Few-shot Synthetic Code
Exact (Baseline) 45.71 45.32 26.69 68.62 59.25 46.17 48.63
PyramidKV 36.80 41.54 18.91 64.88 59.68 42.38 44.03
SnapKV 38.23 42.61 19.07 64.65 59.60 43.28 44.57
StreamingLLM 27.86 39.05 20.70 66.16 58.37 42.91 42.51
Uniform 37.17 42.66 24.31 65.09 58.30 41.40 44.82
BalanceKV 36.97 42.37 23.82 65.49 58.46 42.82 44.99
5 Conclusion
In this paper we present BalanceKV, an algorithm with provable guarantees for reducing the memory complexity of KV cache in modern LLMs. BalanceKV uses a geometric correlated sampling process based on discrepancy theory to recursively compress the KV cache and it provably approximates attention in the streaming setting. We provide empirical evidence to show that our approach leads to lower relative error for single layer attention approximation for various open source LLMs including $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ and $\mathtt{Ministral}$-$\mathtt{8B}$-$\mathtt{Instruct}$-$\mathtt{2410}$, as well as better performance on the LongBench benchmark for long context understanding capabilities of LLMs under various end to end tasks compared to heuristic approaches.

### v2

@@ C.1 Reduction to the INDEX problem
$\mathbb{E}_{x,y\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d}}[\exp(C\langle x,y\rangle/\sqrt{d})]$
$=\left(\frac{1}{2}\exp(Cr^{2}/d^{3/2})+\frac{1}{2}\exp(-Cr^{2}/d^{3/2}))\right)^{d}=\text{cosh}\left(\frac{Cr^{2}}{d^{3/2}}\right)^{d},$
$1\leq\exp(C^{2}r^{4}/4d^{2})\leq\text{cosh}\left(\frac{Cr^{2}}{d^{3/2}}\right)^{d}\leq\exp(C^{2}r^{4}/d^{2})\leq\exp(C^{2})$
where we used the assumption that $r^{2}/d\leq 1$. ∎
Lemma C.4.
Fix $i\in[n]$, select $q_{i}=k_{i}$. Let $|U_{i}|$ be the number of key-value pairs in $U\setminus\{k_{i},v_{i}\}$ whose value vector has non-zero entry in the same coordinate as $v_{i}$.
If $\{k_{i},v_{i}\}\in U$ then with probability $>1-\frac{1}{1000}\cdot\frac{|U_{i}|}{|U|}$
|$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\geq\exp(r^{2}/\sqrt{d})-O\left(\sqrt{|U|}\right).$
Otherwise, with probability $>1-\frac{1}{1000}\cdot\frac{|U_{i}|}{|U|}$
|$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\leq O\left(\sqrt{|U|}\right).$
Proof.
We prove both statements using Chebyshev’s inequality.
In the first case, i.e. when $\{k_{i},v_{i}\}\in U$, the sum contains the term $\exp(\langle k_{i},q_{i}\rangle/\sqrt{d})=\exp(r^{2}/\sqrt{d})$, and otherwise it does not. It therefore remains to upper bound the absolute value of the sum
$X=\sum_{\begin{subarray}{c}\{k,v\}\in U,\\ \{k,v\}\neq\{k_{i},v_{i}\}\end{subarray}}\sigma_{k}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle=\sum_{\begin{subarray}{c}\{k,v\}\in U_{i},\\ \{k,v\}\neq\{k_{i},v_{i}\}\end{subarray}}\sigma_{k}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle,$
$\sigma_{k}\sim\text{Unif}\{-1,1\}$, which effectively introduces “noise” in Bob’s estimate of whether $x_{i}=1$. We upper bound this sum now.
$\mathbb{E}[X]=0$, and $Var(X)=|U_{i}|\cdot Var_{x,y\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d}}(\exp(\langle x,y\rangle/\sqrt{d}))\leq O\left(|U_{i}|\right)$ because
$\begin{split}Var_{x,y\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d}}(\exp(\langle x,y\rangle/\sqrt{d}))\leq\mathbb{E}_{x,y\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d}}[\exp(2\langle x,y\rangle/\sqrt{d})]\leq\exp(4),\end{split}$
by Lemma C.3. We therefore get by Chebyshev’s inequality
$\Pr\left[|X|\geq 1000\sqrt{|U|}\right]\leq\frac{1}{1000}\cdot\frac{|U_{i}|}{|U|}.$
Therefore, with probability $1-\frac{1}{1000}\cdot\frac{|U_{i}|}{|U|}$
|$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\geq\exp(r^{2}/\sqrt{d})-1000\sqrt{|U|}.$
In the second case, the entire sum equals $X=\sum_{k\in U_{i}}\sigma_{k}\exp(\langle k,q_{i}\rangle/\sqrt{d})$. As shown above, $\Pr\left[|X|\geq 1000\sqrt{|U|}\right]\leq\frac{1}{1000}\cdot\frac{|U_{i}|}{|U|}$. Hence, with probability $1-\frac{1}{1000}\cdot\frac{|U_{i}|}{|U|}$
|$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\leq 1000\sqrt{|U|}.$
Corollary C.5.
Suppose bits $i_{1},\ldots,i_{d}$ form a group  – that is, $\pi(i_{1})_{1}=\pi(i_{2})_{1}=\ldots=\pi(i_{d})_{1}$. Then all $v_{i_{1}},\ldots,v_{i_{d}}$ have different non-zero coordinates, and therefore $\sum_{j=1}^{d}|U_{i_{j}}|\leq|U|$.
Therefore, by the union bound argument, the conclusion of Lemma C.4 holds for all $d$ bits which form one group simultaneously with probability 0.999.
Lemma C.6.
Fix a bit $i$. With probability 0.98 the following holds:
  1. 1.
    The error of the approximating algorithm in the only non-zero coordinate of $v_{i}$ is bounded by
    $O\left(\frac{\varepsilon}{\sqrt{d}}\cdot\|\text{softmax}(K\cdot q)\|_{2}\cdot\|V\|_{F}\right).$
  2. 2.
    If the $i$-th bit is 1 then
    |$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\geq\exp(r^{2}/\sqrt{d})-O\left(\frac{\sqrt{|U|}}{\sqrt{d}}\right),$
    and
    |$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\leq O\left(\frac{\sqrt{|U|}}{\sqrt{d}}\right).$
    otherwise.
Proof.
We may think that the process of generating the dataset and the approximating streaming algorithm has the following order: first Alice and Bob jointly generate the partition $\pi$, the key vectors $\tilde{k}_{1},\ldots,\tilde{k}_{n/d}$ and the value vectors $v_{1},\ldots,v_{n}$ all using public randomness. To generate Bob’s input position $i\in[n]$ we generate pair $a\sim\text{Unif}[n/d]$, $b\sim\text{Unif}[d]$ and declare $i=\pi^{-1}(a,b)$. We may assume that $a$ is chosen before the datasets $K$ and $V$ are generated, and $b$ – after.
Before $b$ is drawn, the key vector $k_{i}$ of $i=\pi^{-1}(a,b)$ is already defined, as well as the datasets $K$, $V$ and $U$. Alice can therefore already apply the streaming algorithm to $U$, and Bob can already apply it to $q_{i}=k_{i}$. Therefore, the error vector which the streaming algorithm yields when applied to $q_{i}=k_{i}$ is also defined before $b$ is known.
Clearly, there are no more than $0.0001\cdot d$ coordinates in which the error of approximation exceeds $10000\cdot\frac{\varepsilon}{\sqrt{d}}\cdot\|\text{softmax}(K\cdot q)\|_{2}\cdot\|V\|_{F}.$ Since every value vector has only one non-zero entry, there are no more than $0.0001\cdot d$ coordinates where at least $10000\cdot\frac{U}{d}$ of value vectors from $U$ have non-zero value. We call all coordinates which are in neither of these two groups safe. From the above, at least $99\%$ of the coordinates are safe. Recall that $b\sim\text{Unif}[d]$, and choosing $b$ is equivalent to choosing the coordinate in which $v_{i}$ is non-zero. Therefore, with probability 0.99 over the choice of $b$ the only non-zero coordinate of $v_{i}$ is safe.
At the same time, similarly to Lemma C.4, by Chebyshev inequality, if $U_{i}\subset U$ is the set of all key-value pairs in $U$ whose value vector has the same non-zero coordinate as $v_{i}$ then with probability 0.999 if the $i$-th bit is 1 then
|$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\geq\exp(r^{2}/\sqrt{d})-O\left(\sqrt{|U_{i}|}\right),$
and
|$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\leq O\left(\sqrt{|U_{i}|}\right).$
otherwise.
By union bounding over these two events, we get that the statement of the lemma is correct with high constant probability.
Conclusion of the proof. Let $U^{i}\subset U$ be the set of all pairs from $U$ with the same key as $\{k_{i},v_{i}\}$. Since Alice’s string is drawn from $\text{Unif}\{-1,1\}^{n}$, with probability 0.9 $|U^{i}|\geq 0.4\cdot d$. This is because every bit in the same group as $k_{i}$ belongs to $U$ with probability 1/2.
Observe that by Chebyshev inequality, with high probability 0.999, the denominator of $\text{softmax}(K\cdot q_{i})$ lies in range
$\left[|U^{i}|\cdot\exp(r^{2}/\sqrt{d})+\frac{1}{5}\cdot|U|,|U^{i}|\cdot\exp(r^{2}/\sqrt{d})+20\cdot|U|\right],$
This is because every summand in the denominator, except for $\exp(\langle k_{i},q_{i}\rangle/\sqrt{d})$, is distributed as $\exp(\langle x,y\rangle/\sqrt{d})$, $x,y\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d}$, and the expectation and the variance of this distribution, as shown in Lemma C.3, is $\Theta(1)$. This range is contained in $\left[\frac{1}{5}\cdot(\max\{de^{r^{2}/\sqrt{d}},|U|\}),20\cdot(\max\{de^{r^{2}/\sqrt{d}},|U|\})\right]$. We will denote the denominator as $D$.
Similarly, by Chebyshev inequality, with probability 0.999 the numerator of $\text{softmax}(K\cdot q_{i})$ lies in
$\left[\sqrt{|U^{i}|\cdot\exp(2r^{2}/\sqrt{d})+\frac{1}{5}\cdot|U|},\sqrt{|U^{i}|\cdot\exp(2r^{2}/\sqrt{d})+200\cdot|U|}\right]$
which, since $|U|\leq d\exp(2r^{2}/\sqrt{d})$, is bounded by $\sqrt{200}\cdot\sqrt{d}e^{r^{2}/\sqrt{d}}$.
Suppose that the $i$-th bit is 1. Then,
  • When $\frac{1}{\varepsilon}\geq\sqrt{d}e^{r^{2}/\sqrt{d}}$, by selecting $|U|=\frac{c}{C}\cdot de^{2r^{2}/\sqrt{d}}$ for some enough constant $C$ the value written in the only non-zero coordinate of $v_{i}$ is at least $\frac{e^{r^{2}/\sqrt{d}}}{D}-\frac{1}{100}\frac{e^{r^{2}/\sqrt{d}}}{D}$ and at most $\frac{1}{100}\frac{e^{r^{2}/\sqrt{d}}}{D}$ otherwise, as follows from Lemma C.4;
  • When $\frac{1}{\varepsilon}<\sqrt{d}e^{r^{2}/\sqrt{d}}$, by selecting $|U|=\frac{c}{C}\cdot\frac{1}{\varepsilon^{2}}$ for some big enough constant $C$ the value written in the only non-zero coordinate of $v_{i}$ is at least $\frac{e^{r^{2}/\sqrt{d}}}{D}-\frac{1}{100}\cdot\frac{1}{\varepsilon\cdot D\cdot\sqrt{d}}$, and at most $\frac{1}{100}\cdot\frac{1}{\varepsilon\cdot D\cdot\sqrt{d}}$ otherwise, as follows from Lemma C.4.
The error which the approximator can have in the non-zero coordinate of $v_{i}$ is bounded by
$10000\frac{\varepsilon}{\sqrt{d}}\cdot\|\text{softmax}(K\cdot q)\|_{2}\cdot\|V\|_{F}\leq 10000\frac{\varepsilon}{\sqrt{d}}\cdot\frac{\sqrt{20}\cdot\sqrt{d}e^{r^{2}/\sqrt{d}}}{D}\cdot\sqrt{|U|},$
as shown in Lemma C.6. Below, we show that this error is smaller than the gap between $\frac{1}{40}\cdot\frac{e^{r^{2}/\sqrt{d}}}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$ and the value written in the coordinate, which means that, even though the approximator introduces some error, Bob is still capable to tell whether the $i$-th bit is 1 or 0.
  • When $\frac{1}{\varepsilon}\geq\sqrt{d}e^{r^{2}/\sqrt{d}}$, the gap between the value written in the coordinate and $\frac{1}{40}\frac{e^{r^{2}/\sqrt{d}}}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$ is at least $\frac{1}{1000}\cdot\frac{e^{r^{2}/\sqrt{d}}}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$, and the error
    $\frac{10000\varepsilon}{\sqrt{d}}\cdot\frac{\sqrt{20}\cdot\sqrt{d}e^{r^{2}/\sqrt{d}}}{D}\cdot\sqrt{|U|}\leq\frac{\varepsilon}{2000}\cdot\frac{\sqrt{d}\exp(2r^{2}/\sqrt{d})}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}\leq\frac{1}{2000}\cdot\frac{e^{r^{2}/\sqrt{d}}}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$
    by an appropriate choice of $C$.
  • When $\frac{1}{\varepsilon}<\sqrt{d}e^{r^{2}/\sqrt{d}}$, the gap between the value written in the coordinate and $\frac{1}{40}\frac{e^{r^{2}/\sqrt{d}}}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$ is at least $\frac{1}{1000}\cdot\frac{e^{r^{2}/\sqrt{d}}}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$ and the error
    $\frac{10000\varepsilon}{\sqrt{d}}\cdot\frac{\sqrt{20}\cdot\sqrt{d}e^{r^{2}/\sqrt{d}}}{D}\cdot\sqrt{|U|}\leq\frac{\varepsilon}{2000}\cdot\frac{\exp(r^{2}/\sqrt{d})}{\max\{de^{r^{2}/\sqrt{d}},|U|\}\cdot\varepsilon}\leq\frac{1}{2000}\cdot\frac{\exp(r^{2}/\sqrt{d})}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$
    by an appropriate choice of $C$.